OpenStudy (anonymous):

The sum of the digits of a three digit number is 15. The TENS PLACE digit is TWICE the HUNDREDS PLACE digit. The ONES DIGIT is one less than the hundreds place digit. Find the three digit number. Solve using a SYSTEM OF EQUATIONS.

OpenStudy (anonymous):

x+y+z=15 y=2x z=x-1 x+y+z=15 2x-y=0 x-z=1 can yu go from there?

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (anonymous):

Actually, it asks for three equations and three unknowns.

OpenStudy (anonymous):

well. those are your 3 equations and your 3 unknowns are x, y, z